OpenStudy (anonymous):

What is the solution of log3x −725 = 2? x = −2 x = 2 x = 3 x = 4

OpenStudy (♪chibiterasu):

Are you sure it;s 7^25?? o-o

OpenStudy (cwrw238):

thats a huge number!

OpenStudy (anonymous):

It is log 3x-7 those are supposed to be subscript and then the 25 is a whole # and then says =27

OpenStudy (campbell_st):

what is the base of the log.... and is the question log(3x - 725) = 2

OpenStudy (anonymous):

I can't be log(3x) -725 = 2

OpenStudy (anonymous):

log(3x - 725) = 2 is possible yes

OpenStudy (anonymous):

its a small 7 normal 25 not 725 ok

OpenStudy (anonymous):

I guess the base is 10

OpenStudy (anonymous):

small 7 ?

OpenStudy (campbell_st):

ok... so the base of the log is (3x-7) and the number is 25 so you are dealing with a base 5 log since 25 = 5^2 and the base of 5 wll result in an answer of 2 so you are solving 3x - 7 = 5 so just solve for x

OpenStudy (campbell_st):

and the answer is in the list of choices.

OpenStudy (campbell_st):

this is why \[\log_{5}(25) = \log_{5}(5^2) = 2\]

OpenStudy (campbell_st):

I'm nor sure if Latex is working

OpenStudy (anonymous):

It's no :P

OpenStudy (anonymous):

not*

OpenStudy (anonymous):

Really a problem on a math site

OpenStudy (campbell_st):

ok.. so reading the information the base of the log is 3x - 7 to take the log of 25 and get an answer of 2 then base has to be 5 so as I said before, solve 3x - 7 = 5

OpenStudy (anonymous):

can you show what happened to the 27 on the other side of the equal sign

OpenStudy (anonymous):

The value of a dirt bike decreases by 20% each year. If you purchased this dirt bike today for $500, to the nearest dollar how much would the bike be worth 5 years later? $84 $119 $164 $222

OpenStudy (campbell_st):

here is the solution and reasoning

OpenStudy (campbell_st):

and 27 doesn't exist in the question...

OpenStudy (campbell_st):

and the other question simply requires you to use the compound interest formula.. good luck